OpenStudy (anonymous):

Which of the following is the most simplified version of 48x^5 sqrt ?

OpenStudy (solomonzelman):

you mean \(\normalsize\color{blue}{ \sqrt{48x^5}}\) ?

OpenStudy (anonymous):

yea

OpenStudy (solomonzelman):

Yes, so we can simplify this the following way, \(\normalsize\color{black}{ \sqrt{48x^5}=}\) \(\normalsize\color{black}{ \sqrt{16\times 3 \times x^4\times x}=}\) \(\normalsize\color{black}{ \sqrt{4^2\times 3 \times (x^2)^2\times x}=}\) \(\normalsize\color{black}{ 4x^2\sqrt{3 \times x}=}\) \(\normalsize\color{black}{ 4x^2\sqrt{3 x}.}\)

OpenStudy (solomonzelman):

makes sense ?

OpenStudy (anonymous):

thxs so much

OpenStudy (anonymous):

and yea it does